Please help me solve this equation.
3x/x+2+7/x-2=3

Answer by ikdeep(226)

3x/x+2 +7/x-2=3
take the LCM of the denominator i.e (x+2)(x-2)....and we get
[3x(x-2) + 7(x+2)]/(x+2)(x-2) = 3
In numerator we simply multiplyand we get.........
[3x^2 -6x + 7x + 14]/(x+2)(x-2) = 3
now, in denominator, we applies the identity (a+b)(a-b)= a^2 - b^2...and we get.
[3x^2 -6x + 7x + 14]/ (x)^2 - (2)^2 = 3
[3x^2 -6x + 7x + 14]/ x^2 - 4 = 3
now on cross multipication we get....

3x^2 -6x + 7x + 14= 3(x^2 - 4 )
3x^2 + x + 14= 3x^2 - 12
x + 14 = -12 (3x^2 from both side would be cancelled).
or x = -12 -14
x = -26
,,,,,,Now to verify our answer,,put the value x = -26,,,in the given equation and you will certainly get "3" as answer which is equal to RHS i.e "3".
Please feel free to revert back for any further queries.